NOAA-17, Level 1B, Data Availability for Testing, 15 July 2002 - Issued: 15 July 2002; 1736 UTC

NOAA-17, Level 1B, Status of NOAA-17 Level 1B Data Availability.

Date/Time(UTC) Message Issued:
15 July 2002 1736 UTC

Satellite(s) Involved:

Instrument(s) Involved:

Product(s) Involved/Affected:
Level 1B

Date/Time(UTC) of Initial Implementation:
15 July 2002

Details/Specifics of Change:
NOAA-17 was handed over from NASA to NOAA on July 14. SOCC has issued a
series of NOAA-17 status reports, the latest of which is provided below.
The Information Processing Division is making the level 1B and 1B* data
available for user check out as we continue our verification and validation
phase. Please note that this data has not been declared operational and
changes may be made to the data as anomalies are identified and corrected.
The user community will be notified of changes but not in accordance with
the user notification policy. The policy will be enforced once the data is
declared operational. The following anomalies have been identified and
corrected where possible:

- Advanced Microwave Sounding Unit-A1-1 - An intermittent timing error was
present in the data stream. On July 9, SOCC was able to correct the problem
by turning the AMSU-A1-1 instrument off and back on (2241Z to 2244Z).

- HIRS - On July 11, as of orbit B0023637, the scan mirror temperature
sensor failed completely. Prior to this, it was intermittent. This should
have no impact on level 1B and product processing. Michael Chalfant, Office
of Research and Application, examined imagery produced from the NOAA-17 HIRS
data and made recommendations for updates to the level 1B processing. On
the afternoon of July 11, the following changes were introduced into
1. Due to possible contamination in the space view samples, the number of
space view samples used in processing has been reduced from 45 to 40 making
it compatible with NOAA-16.
2. A constant roll type attitude correction of 1.8 degrees was introduced
to correct an apparent across track error that appears to be the same as in

SOCC status report:
NOAA-17 Status as of 7/15/02
The NOAA-M spacecraft was launched Monday June 24th at 2:23pm Eastern Day
light savings time -175/18:23z.

All deployments were successful.

All of the instruments have been turned on:
AVHRR VIS CH (1,2,3A) 6/24/02 175/2115z
AMSU-B Turn on 6/25/02 176/1840z
SARR Turn on 6/27/02 178/2252z
SARP Turn on 6/28/02 179/1512z
AMSU-A1 Turn on 6/28/02 179/1830z
AMSU-A2 Turn on 6/28/02 179/2010z
SBUV Turn on 6/30/02 181/1430z
DCS Turn on 7/01/02 182/1400z
AVHRR (IR) Turn on 7/10/02 191/1545z
HIRS Turn on 7/10/02 191/1900z
SEM Turn on 7/11/02 192/1520z

Handed over NOAA-17 from NASA to NOAA on Sunday July 14th.
Completed the on-line OV tests for SEM
Continued to investigate the HIRS Scan Mirror temp flag red high.

Investigating incorrect data causing red flag since other temperatures
around area appear
to be nominal.
Continued to monitor Thruster 6 temp.

Investigation indicates a faulty sensor relaying incorrect data.
All other temperatures in the surrounding areas are nominal.

Other Thruster temperatures running between 14 and 18 degrees
Continued to monitor AMSU-A1

Have not seen the data dropouts since the instrument was cycled off/on
on 7/9/02
Tuesday at 6:45pm

We are investigating a space view/warm view bias in channel 3
We will continue to monitor the AMSU-A1 data.

No problems

as of 196/1400z (7/15/02 10:00am)
5 OV tests were scheduled
5 OV tests were completed

To Date:
95 OV tests scheduled
96 OV tests completed ( additional off-line data analysis continues on
some tests)

Begin the space view testing on AMSUA1 & A2
Complete the on-line OV testing on HIRS
Continue OV tests on SBUV

Contact Person(s) Name/Email/Phone Number for Questions:
Emily D. Harrod/ ext 117

Web site(s) that contain other relevant information (where applicable): - SOCC morning reports - History of level 1B